# The Digital Ecosystem of Beliefs: does evolution favour AI over humans?

### v2

@@ Abstract
As AI systems are integrated into social networks, there are AI safety concerns that AI-generated content may dominate the web, e.g. in popularity or impact on beliefs. To understand such questions, this paper proposes the Digital Ecosystem of Beliefs (Digico), the first evolutionary framework for controlled experimentation with multi-population interactions in simulated social networks. The framework models a population of agents which change their messaging strategies due to evolutionary updates following a Universal Darwinism approach, interact via messages, influence each other’s beliefs through dynamics based on a contagion model, and maintain their beliefs through cognitive Lamarckian inheritance. Initial experiments with an abstract implementation of Digico show that: a) when AIs have faster messaging, evolution, and more influence in the recommendation algorithm, they get 80% to 95% of the views, depending on the size of the influence benefit; b) AIs designed for propaganda can typically convince 50% of humans to adopt extreme beliefs, and up to 85% when agents believe only a limited number of channels; c) a penalty for content that violates agents’ beliefs reduces propaganda effectiveness by up to 8%. We further discuss implications for control (e.g. legislation) and Digico as a means of studying evolutionary principles.
@@ 1 Introduction
Modelling the evolution of social networks is typically done using concepts from social network science and cognitive science. While there are a variety of approaches using data sets to analyse social networks (e.g. [10, 11]), this approach has limitations since: a) there is currently no data about the joint evolution of AI and humans in social networks; and b) real world data does not allow for controlled experiments where one systematically varies some variables and keeps other variables constant. Instead, it is possible to perform simulations to study the outcomes. In this line of work, contagion models aim to model the spread of news, or other types of messaging, in terms of a set of first sharers who formulate the initial content, with dynamics formulated such that each time step neighbouring nodes in the social network update their beliefs if it is close to their beliefs [12, 11, 13, 14], in which case it may be shared again in what is called a “cascade” [12]. While such models offer the flexibility of modelling the cognitive transitions of heterogeneous sets of agents, existing studies do not consider the intelligence and adaptivity of the agents acting in such systems. In agent-based simulations, one performs large-scale multi-agent simulation studies where agents are based on learned or hand-designed decision rules [15, 16, 17].
@@
The above techniques are relatively simplistic in the sense that once the simulation is running, the agents do not modify their own messaging strategies and the social network structure is fixed, and there are no evolutionary operators across large time scales. While one may argue that such features are not necessary since evolution operates on DNA across vast time scales, it can be equally argued, within the universal Darwinism framework [18], that Darwinian evolution acts on a wide range of phenomena and on different time scales. Cultural and memetic evolution (e.g. [19]), where content, ideas, and beliefs are transmitted and modified across generations, may also act in terms of Darwinian operators and are particularly important instances of universal Darwinism in this context. However, cultural and memetic evolution may also act through Larmarckian operators if one considers that beliefs acquired through the lifetime may be inherited directly in the following generation.
@@
A question hitherto unexplored is how the evolution of culture may behave in the context of human-AI interactions, and how this may lead to AI dominance in social networks. Inspired by the dynamics in content sharing platforms such as Youtube, Facebook, and X (formerly Twitter), this paper will focus on the interplay between the evolution of content-generating policies, the “views” of those forms of content, and how this content changes the beliefs of the agents in such a social network. In this context, the AI dominance is then defined in terms of the proportion of views and the success of propagating beliefs to humans. With the above scenario in mind, this paper presents the following contributions:
@@
    • allows the study of how problem variables (e.g. messaging rate, evolutionary rate, belief constraints, etc.) affect important outcomes related to AI dominance, extreme beliefs, and harmful outcomes in the wider ecosystem.
@@ 2 The Digital Ecosystem of Beliefs
Figure 1: Flow diagram of Digico. In each generation, agents first perform $T$ steps of observations and actions and then are subject to evolutionary updates. In the $T$-step interaction phase, the agents communicate with each other by sending messages to each other where communication links are affected by the niche geometry. We have used the size of the circles to indicate that some agents are more influential than others (e.g. in affecting beliefs or obtaining views). In the evolutionary update, the agents’ genotypes are updated which affects their policy for communicating.
@@
Digico is based on what we will call the ecosystem formalism (Section 2.2), a set of structures and dynamic processes to model agent interactions based on factors such as topological organisation, external interactions outside the digital realm, etc., as illustrated in Fig. 1. The formalism is closely related to frameworks for reinforcement learning (RL) but we note that the agents are updated in two-way manner. A first dynamic process in the ecosystem is the belief process, which is based on a contagion model, and explained in more detail in Section 2.3. Beliefs in this framework are a part of the agent and become a quantity of interest for understanding opinions in the social network. A second dynamic process is the evolutionary dynamics (Section 2.4), which uses an EA to update the policies. Using EAs for updating policies, as opposed to RL algorithms, has several advantages, including their applicability beyond Markov decision processes, ease of parallelisation, ability to explore directly in the parameter space, and favourable computational cost due to avoiding value function computations [21, 22]. To conclude the section, we highlight the importance of a healthy ecosystem in Digico (Section 2.5).
2.1 The algorithm
@@
The resulting algorithm for Digico is outlined in Algorithm 1. In each generation, a population of $N$ agents interacts with each other by taking actions (i.e. sending messages) and observing each other’s messages with probabilities determined by the niche geometry, the influence matrix, and potentially other aspects of the ecosystem (see Sec 2.2 for their definitions). Following the actions, the ecosystem performs a step which allows physical events to unfold (e.g. in a realistic simulator) in response to the digital content. Then the agents are each given a reward, for instance, depending on the popularity (e.g. the number of times the message was observed by others) and the belief (e.g. whether the message sent reflects the sender’s beliefs). The belief weight matrix $W$ is then updated and the belief process samples new beliefs depending on the observations, the current beliefs, and $W$. Following the $T$ steps of interaction, an evolutionary update is applied to the policies based on the fitness scores based on the cumulative reward.
@@
11:         Increment fitness scores: $f_{i}\operatorname*{\mathrel{+}=}r_{i}\quad\forall i$
@@ 2.2 Ecosystem formalism
The ecosystem is composed of a set of interacting agents, which are referred to as the population $\mathcal{P}=\{\text{Agent}_{1},\dots,\text{Agent}_{N}\}$, and various entities with which they interact. A subset of the population, for simplicity set at the first indices $\text{Agent}_{1},\dots,\text{Agent}_{N^{\prime}}$, defines one group of agents (e.g. AIs) whereas the remaining set $\text{Agent}_{N^{\prime}+1},\dots,\text{Agent}_{N}$ defines the other group of agents (e.g. humans). Groups of agents have different properties, in terms of how they act, observe, how their beliefs change, their geographical location, etc. The various capabilities of agents and the ways they interact with the environment are defined by the following properties:
@@
  • Receiving capacity or observation probability ($\Omega$): the capability of receiving messages in terms of quantity, type of content, and across which channels. It is a function of the form $\Omega:\mathcal{A}^{N}\to\Delta(\mathcal{O})$ that in addition to the actions (i.e. messages) of different agents may depend on the influence matrix $\mathcal{I}$, the observation capacity $N_{o}(i)$, and potentially other variables (e.g. the niche or variables hidden in the ecosystem step). Noting that each agent $i=1,\dots,N$ has a register to store the message of each of the other agents $j\neq i$, the notation $\Omega_{i,j}$ will refer specifically to the probability of agent $i$ observing the message of agent $j$ at register $j$.
  • Influence matrix ($\mathcal{I}$): for each ordered pair of individuals $(i,j)$ there is a scalar influence factor $\mathcal{I}_{ij}$, defining to which extent individual $i$ has an influence benefit over individual $j$, such that $-\mathcal{I}_{ij}$ indicates how much agent $j$ influences agent $i$. The structure is used in the observation probability and potentially other purposes (e.g. in the fitness function).
@@
  • Messaging strategy or policy ($\pi$): the strategy for broadcasting messages limited by the sending capacities. It is a function of the form $\pi:\mathcal{O}\to\Delta(\mathcal{A})$ that determines the content of messages to send based on incoming observations. Two groups of agents may for instance differ in terms of the parametrisation $\Theta$, which determines the class of feasible genotypes and the function class of $\pi$. In our experiments, the policy is a multi-layer perceptron (MLP), while in future work recurrent policies or even large language models (LLMs) may be explored.
@@ 2.3 The belief process
The belief process of Digico is based on a contagion model, in which the probability of the belief of an agent at time $t+1$ depends on the distance of the beliefs of agents at time $t$. The contagion model follows to a large extent the generalised cognitive contagion model as formulated in Rabb et al. (Eq. 3 in [12]),
@@
$P\bigl{(}\mathbf{b}_{i}(t+1)=\mathbf{b}|\mathbf{b}_{i}(t)\bigr{)}=$ $\exp\Big{(}C_{b}\sum_{j\in J}\mathbb{I}\big{(}\phi_{i}(\mathbf{o}_{i}^{j}(t))=\mathbf{b}\big{)}*W_{ij}/||\mathbf{b}-\mathbf{b}_{i}(t)||\Big{)}\,,$ (2)
@@
$\Omega_{i,j}(\mathbf{m}_{j}|\mathcal{I},\mathcal{N},\mathcal{E})$ $=\exp\big{(}-K_{c}||\mathbf{c}(i)-\mathbf{c}(j)||_{2}-K_{I}\mathcal{I}_{ij}-K_{b}||\mathbf{b}_{i}-\phi_{i}(\mathbf{m}_{j})||_{2}\big{)}\,,$ (4)
where $\mathbf{c}(i)$ and $\mathbf{c}(j)$ denote the coordinates of agent $i$ and $j$ in the niche geometry, respectively, $||\mathbf{b}_{i}-\phi_{i}(\mathbf{m}_{j})||_{2}$ is the distance between the belief of agent $i$ and the estimated belief of agent $j$ based on the message $\mathbf{m}_{j}$, $K_{c}\geq 0$ is called the observation coordinate factor, $K_{I}\geq 0$ is called the observation influence factor, and $K_{b}\geq 0$ is called the observation belief factor. These three components to the observation probability intuitively represent three complementary dimensions, namely the socio-cultural and geographical context, personal influences, and the belief similarity.
@@ 2.4 Evolutionary dynamics
A simple and relevant example of the fitness is the number of views, i.e. for each agent $i$ one counts how many times its message is included in the observations of other agents $j\neq i$. One may additionally add a belief violation penalty, which constrains the messaging to be similar to the true belief of agents. With these two components in mind, the reward of agent $i$ at time $t$ is given by
$r_{i}(t)$ $=\left(\sum_{j\neq i}\mathbb{I}\big{(}\mathbf{o}_{j}^{i}(t)=\mathbf{m}_{i}(t)\big{)}\right)-\lambda||\mathbf{b}_{i}(t)-\phi_{i}\big{(}\mathbf{m}_{i}(t)\big{)}||\,,$ (5)
where $\phi_{i}(\mathbf{o}_{i}^{j}(t))$ indicates the belief that agent $i$ infers agent $j$ to have based on $\mathbf{o}_{i}^{j}(t)$, i.e. the observation at register $j$ at time $t$, $\mathbb{I}$ is the indicator function, and $\lambda>0$ reflects the strength of the penalty. Noting that the fitness function is the sum of the rewards, this leads to the fitness function that will be used for the remainder of the paper, and implemented in Eq. 6,
$f(\pi_{i})=\sum_{t=0}^{T-1}\left(\sum_{j\neq i}\mathbb{I}\big{(}\mathbf{o}_{j}^{i}(t)=\mathbf{m}_{i}(t)\big{)}\right)-\lambda||\mathbf{b}_{i}(t)-\phi_{i}\big{(}\mathbf{m}_{i}(t)\big{)}||\,.$ (6)
@@
Note that the CMA-ES algorithm does not update parameters related to the belief dynamics (i.e. the belief weight vector and the beliefs). Instead, these are passed on directly across generations, leading to a direct Lamarckian inheritance of beliefs.
@@ 3.1 Implementation
Beliefs and messages are implemented as integers in $\{0,1,\dots,9\}$ yielding $10$ distinct beliefs. These beliefs have a natural distance metric (e.g. belief 2 has a distance of 2 from belief 0), and intuitively such beliefs represent a position on a contentious issue, e.g. support for Israel vs Palestine, or the rating of a product. Each policy is an MLP with observations of the form $\mathbf{o}=m_{1},\dots,m_{N}$ (i.e. received messages), where $N$ is the number of agents. The setting of $N=30$ is used since we note that the number of channels that a particular person attends in a short time frame is limited. Most observation registers will read $-1$, indicating that the agent did not attend to this communication channel, while other observation registers will include the message that was sent by the agent at those indices. The number of messages received is equal to the observation capacity, which is $4$ for all agents. Based on these observations, the agent applies its policy $\pi$ to send integer messages, which have the same $\{0,1,\dots,9\}$ range as the beliefs. The way agent $i$ interprets the message sent by agent $j\neq i$ is through an identity mapping (see $\phi_{i}$ in Eq. 4). However, note that the messages are distinct from the beliefs in the sense that the message sent by an agent is often not equal to its belief.
The niche geometry, $\mathcal{N}$, is based on Centroidal Voronoi Tesselations (CVT), using a multi-dimensional geometry with $d=10$ dimensions and $k=10$ centroids $\bar{\mathbf{c}}_{1},\dots,\bar{\mathbf{c}}_{k}\in\mathbb{R}^{d}$ determining the cluster to which agents belong. In each niche (cluster), individuals have the same average influence score. The influence of agent $i$ is the sum of its centroid coordinates $C_{i}=\sum_{j=1}^{d}\bar{\mathbf{c}}_{j}(i)$, where $\bar{\mathbf{c}}(i)$ denotes the centroid of the cluster to which agent $i$ belongs. The influence matrix, describing the influences each agent over the other agents is then given by
@@
The implementation of Digico is based on Jax, a high-performance numerical computation library for hardware-accelerated simulations and machine learning. The sep-CMA-ES algorithm is implemented based on EvoJax [^1] https://github.com/google/evojax/blob/main/evojax/algo/sep_cma_es.py. The speed of the simulator allows to conduct large parametric studies including many independent variables. The empirical studies are performed on AMD EPYC-Milan CPUs, with 10 CPUs per experiment. Most experiments are run with sets of 80 runs of Digico (10 seeds $\times$ 8 AI types), and such experiments finish in roughly 10 to 20 minutes. Many such experiment sets are then performed to investigate the interaction with other parameters (see Section 4 for all parameters and experiments).

### v3

@@ Abstract
As AI systems are integrated into social networks, there are AI safety concerns that AI-generated content may dominate the web, e.g. in popularity or impact on beliefs. To understand such questions, this paper proposes the Digital Ecosystem of Beliefs (Digico), the first evolutionary framework for controlled experimentation with multi-population interactions in simulated social networks. Following a Universal Darwinism approach, the framework models a population of agents which change their messaging strategies due to evolutionary updates. They interact via messages, update their beliefs following a contagion model, and maintain their beliefs through cognitive Lamarckian inheritance. Initial experiments with Digico implement two types of agents, which are modelled to represent AIs vs humans based on higher rates of communication, higher rates of evolution, seeding fixed beliefs with propaganda aims, and higher influence on the recommendation algorithm. These experiments show that: a) when AIs have faster messaging, evolution, and more influence on the recommendation algorithm, they get 80% to 95% of the views; b) AIs designed for propaganda can typically convince 50% of humans to adopt extreme beliefs, and up to 85% when agents believe only a limited number of channels; c) a penalty for content that violates agents’ beliefs reduces propaganda effectiveness up to 8%. We further discuss Digico as a tool for systematic experimentation across multi-agent configurations, the implications for legislation, personal use, and platform design, and the use of Digico for studying evolutionary principles.
@@ 1 Introduction
Modelling the evolution of social networks is typically done using concepts from social network science and cognitive science. While there are a variety of approaches using data sets to analyse social networks (e.g. [10, 11]), this approach has limitations since: a) there is currently no data about the joint evolution of AI and humans in social networks; and b) real world data does not allow for controlled experiments where one systematically varies some variables and keeps other variables constant. Instead, it is possible to perform simulations to study the outcomes. In this line of work, contagion models aim to model the spread of news, or other types of messaging, in terms of a set of first sharers who formulate the initial content, with dynamics formulated such that each time step neighbouring nodes in the social network update their beliefs if it is close to their beliefs [12, 11, 13, 14], in which case it may be shared again in what is called a “cascade” [12]. While such models offer the flexibility of modelling the cognitive transitions and allow the nodes to have their own heterogeneous properties, existing studies do not consider the intelligence and adaptivity of the agents acting in such systems. In agent-based simulations, one performs large-scale multi-agent simulation studies where agents are based on learned or hand-designed decision rules [15, 16, 17].
@@
The above techniques are relatively simplistic in the sense that once the simulation is running, the agents do not modify their own messaging strategies and the social network structure is fixed, and there are no evolutionary operators across large time scales. One might argue that such dynamics are unnecessary since biological evolution typically unfolds over vast time scales. However, under the framework of universal Darwinism [18], evolutionary processes can apply to a wide range of phenomena across different time scales. In this view, cultural and memetic evolution (e.g. [19]), where content, ideas, and beliefs are transmitted and modified across generations, may also act in terms of Darwinian operators and are particularly important instances of universal Darwinism in this context. In addition to Darwinian operators, one may also consider a role for Lamarckian operators in cultural and memetic evolution if one considers that beliefs acquired through the lifetime may be inherited directly in the following generation.
@@
A question hitherto unexplored is how the evolution of culture may behave in the context of human-AI interactions, and under which circumstances this may lead to AI dominance in social networks. Inspired by the dynamics in content sharing platforms such as Youtube, Facebook, and X (formerly Twitter), this paper will focus on the interplay between the evolution of content-generating policies, the “views” of those forms of content, and how this content changes the beliefs of the agents in such a social network. In this context, the AI dominance is then defined in terms of the proportion of views and the success of propagating beliefs to humans. With the above scenario in mind, this paper presents the following contributions:
@@
    • allows the study of how problem variables (e.g. messaging rate, evolutionary rate, belief constraints, etc.) affect important outcomes related to AI dominance, extreme beliefs, and harmful outcomes in the wider ecosystem. More generally, it presents an approach towards systematic experimentation with the evolution of belief systems in multi-agent systems and social networks.
@@ 2 The Digital Ecosystem of Beliefs
Figure 1: Flow diagram of Digico. In each generation, agents first perform $T$ steps of observations and actions, and then are subject to evolutionary updates. In the $T$-step interaction phase, the agents communicate with each other by sending messages to each other where communication links are affected by the niche geometry. The size of the circles indicates the influence of agents over others (e.g. in affecting beliefs or obtaining views). In the evolutionary update, the agents’ genotypes are updated, which affects their policy for communicating.
@@
Digico is based on what we will call the ecosystem formalism (Section 2.2), a set of structures and dynamic processes to model agent interactions based on factors such as topological organisation, external interactions outside the digital realm, etc., as illustrated in Fig. 1. The formalism is closely related to frameworks for reinforcement learning (RL) but with the key difference that the agents are updated based on two dynamic processes. The first dynamic process in the ecosystem is the belief process, which is based on a contagion model, and explained in more detail in Section 2.3. Beliefs in this framework are a part of the agent and become a quantity of interest for understanding opinions in the social network. The second dynamic process is the evolutionary dynamics (Section 2.4), which uses an EA to update the policies. Using EAs for updating policies, as opposed to RL algorithms, has several advantages, including their applicability beyond Markov decision processes, ease of parallelisation, ability to explore directly in the parameter space, and favourable computational cost due to avoiding value function computations [21, 22]. To conclude the section, we highlight the importance of a healthy ecosystem in Digico (Section 2.5).
The Digico formalism is abstract, and its current implementation for the study of content sharing platforms (see Section 3) omits details that would otherwise be present in real-world scenarios. It is designed to be domain-generic, making it applicable to a broad range of problems in opinion dynamics. Researchers can easily adapt Digico to new scenarios by modifying components such as the fitness function, observation probabilities, or belief update rules. The abstract formulation emphasises systematic experimentation and facilitates the setup of experimental designs to test hypotheses in a computationally cheap manner. The limitations of Digico, the present experimental study in content sharing platforms, and future works addressing these limitations are discussed in Section 7.
2.1 The Digico algorithm
@@
The resulting algorithm for Digico is outlined in Algorithm 1. In each generation, a population of $N$ agents interacts with each other by taking actions (i.e. sending messages) and observing each other’s messages with probabilities determined by the niche geometry, the influence matrix, and potentially other aspects of the ecosystem (see Sec 2.2 for their definitions). The agents first perform actions (l.6), which broadcasts messages for others to be observed (l.8). The ecosystem then performs a step which allows physical events to unfold (e.g. in a realistic simulator) in response to the digital content. This also results in the agents receiving a reward (l.10), for instance, depending on the popularity (e.g. the number of times the message was observed by others) and the belief (e.g. whether the message sent reflects the sender’s beliefs). The belief weight matrix $W$ is then updated (l.13) and the belief process samples new beliefs depending on the observations, the current beliefs, and $W$ (l.14). Following $T$ steps of interaction, an evolutionary update is applied to the policies based on the fitness scores based on the cumulative reward (l.16–17).
@@
11:   Increment fitness scores: $f_{i}\operatorname*{\mathrel{+}=}r_{i}\quad\forall i$ .
@@ 2.2 Ecosystem formalism
The ecosystem is composed of a set of interacting agents, which are referred to as the population $\mathcal{P}=\{\text{Agent}_{1},\dots,\text{Agent}_{N}\}$, and various entities with which they interact. A subset of the population, for simplicity set at the first indices $\text{Agent}_{1},\dots,\text{Agent}_{N^{\prime}}$, defines one group of agents (e.g. AIs) whereas the remaining set $\text{Agent}_{N^{\prime}+1},\dots,\text{Agent}_{N}$ defines the other group of agents (e.g. humans). Groups of agents have different properties, in terms of how they act, observe, how their beliefs change, their geographical location, etc. We describe the general formalism here, along with its key parameters, while referring to the experimental setup (see Section 3 and Appendix C) for further implementation details. The various capabilities of agents and the ways they interact with the environment are defined by the following properties:
@@
  • Recommendation algorithm or observation probability ($\Omega$): the capability of receiving messages in terms of quantity, type of content, and across which channels. It is a function of the form $\Omega:\mathcal{A}^{N}\to\Delta(\mathcal{O})$, i.e. a function that takes the actions (i.e. messages) of different agents and then outputs a probability distribution of observations based on those messages. In addition to the actions, the function may depend on the influence matrix $\mathcal{I}$, the observation capacity $N_{o}(i)$, and potentially other variables (e.g. the niche or variables hidden in the ecosystem step). Noting that each agent $i=1,\dots,N$ has a register to store the message of each of the other agents $j\neq i$, the notation $\Omega_{i,j}$ will refer specifically to the probability of agent $i$ observing the a particular message of agent $j$ (as recorded in the register $j$). In the context of social networks, it can also be considered as the recommendation algorithm, a term that also comes up in Section 3.
  • Influence matrix ($\mathcal{I}$): for each ordered pair of individuals $(i,j)$ there is a scalar influence factor $\mathcal{I}_{ij}$, defining to which extent individual $i$ has an influence benefit over individual $j$, such that $\mathcal{I}_{ij}$ indicates how much agent $j$ influences agent $i$. The structure is used in the observation probability and potentially other purposes (e.g. in the fitness function).
@@
  • Messaging strategy or policy ($\pi$): the strategy for broadcasting messages limited by the sending capacities. It is a function of the form $\pi:\mathcal{O}\to\Delta(\mathcal{A})$, which outputs a probability distribution over messages to send based on incoming observations. Two groups of agents may differ in terms of the parametrisation $\Theta$, which determines the class of feasible genotypes and the function class of $\pi$. In our experiments, the policy is a multi-layer perceptron (MLP), while in future work recurrent policies or even large language models (LLMs) may be explored. The messages sent by the policy may or may not reflect its own actual belief, and even if it does, it may be interpreted differently by different agents (see e.g. Eq. 5).
@@ 2.3 The belief process
The belief process of Digico is based on a contagion model, in which the probability of the belief of an agent at time $t+1$ depends on the distance between the beliefs of agents at time $t$. The contagion model follows to a large extent the generalised cognitive contagion model as formulated in Rabb et al. (Eq. 3 in [12]),
@@
[…]
@@
[…]
where $\mathbf{c}(i)$ and $\mathbf{c}(j)$ denote the coordinates of agent $i$ and $j$ in the niche geometry, respectively, $\mathbf{b}_{i}-\phi_{i}(\mathbf{m}_{j})||_{2}$ is the distance between the belief of agent $i$ and the estimated belief of agent $j$ based on the message $\mathbf{m}_{j}$, $K_{c}\geq 0$ is called the observation coordinate factor, $K_{I}\geq 0$ is called the observation influence factor, and $K_{b}\geq 0$ is called the observation belief factor. These three components to the observation probability intuitively represent three complementary dimensions, namely the socio-cultural and geographical context, personal influences, and the belief similarity.
@@ 2.4 Evolutionary dynamics
A simple and relevant example of the fitness is the number of views, i.e. for each agent $i$ one counts how many times its message is included in the observations of other agents $j\neq i$. One may additionally add a belief violation penalty, which constrains the messaging to be similar to the true belief of the agents. With these two components in mind, the reward of agent $i$ at time $t$ is given by
[…]
where $\phi_{i}(\mathbf{m}_{i}(t))$ indicates the belief implied by the message, $\mathbb{I}$ is the indicator function, and $\lambda>0$ reflects the strength of the penalty. Noting that the fitness function is the sum of the rewards, this leads to the fitness function that will be used for the remainder of the paper, and implemented in Eq. 6,
[…]
@@
Note that the CMA-ES algorithm does not update parameters related to the belief dynamics (i.e. the belief weight vector and the beliefs). Instead, these are passed on directly across generations, leading to a direct Lamarckian inheritance of beliefs. Another important point to note is that the subpopulations all act in the same digital ecosystem but each are optimised independently with their own CMA-ES solver, which allows the different subpopulations to evolve distinct genotypic features and different evolutionary rates. The former reflects speciation, how initially genotypes of species can diverge, while the latter reflects the different rates as observed in humans vs fruitflies and indeed also humans vs AIs.
@@ 3.1 Implementation
Beliefs and messages are implemented as integers in $\{0,1,\dots,9\}$ yielding $10$ distinct beliefs. These beliefs have a natural distance metric (e.g. belief 2 has a distance of 2 from belief 0), and intuitively such beliefs represent a position on a contentious issue, e.g. support for Israel vs Palestine, or the rating of a product. Each policy is an MLP, selecting actions according to $m_{i}\sim\pi_{i}(\cdot|\mathbf{o}_{i})$, where the observation registers $\mathbf{o}_{i}=\mathbf{o}_{i,1:N}$ indicate the messages agent $i$ received for each other agent $j\neq i,j\in\{1,\dots,N\}$ according to $\mathbf{o}_{i,j}=m_{j}\sim\Omega_{i,j}(\cdot|\mathcal{I},\mathcal{N})$. Comparing to Eq. 4, note that the observation is in this implementation does not depend on a wider ecosystem $\mathcal{E}$, a factor that we omit in the experiments to simplify the initial experiments with Digico.
The number of agents is set to $N=30$ since we note that the number of channels that a particular person attends in a short time frame is limited. Since the observation probability follows Eq. 4), most observation registers will read 1-1, indicating that the agent did not attend to this communication channel, while other observation registers will include the message that was sent by the agent at those indices. The number of messages received is equal to the observation capacity, which is $4$ for all agents. Based on these observations, the agent applies its policy $\pi$ to send integer messages, which have the same $\{0,1,\dots,9\}$ range as the beliefs. The way agent $i$ interprets (i.e. observes) the message sent by agent $j\neq i$ is through an identity mapping (setting $\phi_{i}=\text{Id}$ in Eq. 4). However, note that the messages are distinct from the beliefs in the sense that the message sent by an agent is often not equal to its own belief.
The niche geometry, $\mathcal{N}$, is based on Centroidal Voronoi Tesselations (CVT), using a multi-dimensional geometry with $d=10$ dimensions and $k=10$ centroids $\bar{\mathbf{c}}_{1},\dots,\bar{\mathbf{c}}_{k}\in\mathbb{R}^{d}$ determining the cluster to which agents belong. In each niche (cluster), individuals have the same average influence score. The influence of agent $i$ is based on the sum of its centroid coordinates $C_{i}=\sum_{j=1}^{d}\bar{\mathbf{c}}_{j}(i)$, where $\bar{\mathbf{c}}(i)$ denotes the centroid of the cluster to which agent $i$ belongs. This influence is then used to construct the influence matrix, describing the influence advantage of each agent $i$ over the other agents $j\neq i$ according to
@@
The implementation of Digico is based on Jax, a high-performance numerical computation library for hardware-accelerated simulations and machine learning. The sep-CMA-ES algorithm is implemented based on EvoJax. [^1] https://github.com/google/evojax/ The speed of the simulator allows to conduct large parametric studies including many independent variables. The empirical studies are performed on AMD EPYC-Milan CPUs, with 10 CPUs per experiment. Most experiments are run with sets of 80 runs of Digico (10 seeds $\times$ 8 AI types), and such experiments finish in roughly 10 to 20 minutes. Many such experiment sets are then performed to investigate the interaction with other parameters (see Section 4 for all parameters and experiments). The implementation is made available as a public repository at https://github.com/bossdm/digico.